OpenStudy (anonymous):

Solve for c. c plus two-fifths equals negative one-tenth c equals negative one-fifth c equals negative one-half c equals one-tenth c equals three-fifths

OpenStudy (anonymous):

\[c + \left(\begin{matrix}2 \\ 5\end{matrix}\right) = -\left(\begin{matrix}1 \\ 2\end{matrix}\right)\]

OpenStudy (anonymous):

What should you do to get c by itself?

OpenStudy (anonymous):

subtract 2/5 from both sides

OpenStudy (anonymous):

Yes. Do you know how to do that?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Then, that should give you the answer of c.

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

oh well im sorry for waisting your time @algebraprobs